Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day11


Crew 100A Log Book for February 22, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: EVA #10 (14:16 - 14:36), EVA#11 (15:04-15.37), EVA#12 (16:00-17:00), EVA #13 (17:00-18:30)

Brief Narrative of Field Mission Results:

We spent the morning analyzing some of the results of our experiments and writing some of the reports for the two main missions during our stay here in Mars: the Tiger mission and the SALM Mission.

Since we only have 3 more days of research and still a lot to be done, the afternoon started early and we managed to complete 4 EVAs, two as part of the experiment to test the effect of stress induced by communication (both with ATV) and another two to visit the sites requested by the SALM team: P2 (N38.40746º W110.79280º) and P3 (N38.40737º W110.79261º) by foot, and P5 (N38.42638º W110.78342º) and P6 (N38.42638º W110.78342º) by ATV.

It is now time to compile the samples we gathered along with the videos and pictures to send to two teams, one at the La Reunion island (SALM) and another at the ESTEC (ESA) - Netherlands.

Another very busy day that has been rewarded with some nice results.

Engineering/Hab Maintenance: Fixed the GreenHab gray water system (two leaks solved - see engineering report).

Plans for Tomorrow:  Continue with more EVAs to test the influence of stress induced by communication on the EVAs and hopefully sort out the problem with Max.

Reports scheduled for today: Engineering, Commander, Journalist, Science, EVA #10, EVA#11, EVA#12 and EVA#13

Inventory: Received some packages of food (information sent to Jean Hunter). Complete food inventory will be done at the end of the mission.

Support Requested: We only have about 1.5 containers of gas left for the ATVs. Consider bringing some gas during crew change over.

Miscellaneous: N/A


Commander's Report

Ivo Ferreira Reporting

The EVA marathon

Today we almost finished with the habitability experiments, only missing Kévin and Audrey that still have to do the color composition and the copy mirror experiment tomorrow. Then it will be time for Ayako to give us her debriefing. I wonder how close the results from our saliva will be from the questionnaires; they were quite difficult sometimes but "saliva never lies".

Quick lunch and then we went for an afternoon full of EVAs. First it was time for Ayako and Audrey to follow Mission 2 of the communication-stress experiment with me on mission control. ATV problems along with mis-communication made them run out of oxygen before reaching the refill station. May they R.I.P....

Then it was time for Kévin and Romain to test their skills on the same mission. Whatever time they gained on ATV speed they lost due to mis-communication and almost had the same fate as the girls. Luckily their error led then right to the refill station.

It was around 4pm and we still had a lot to do. So, when Kévin and Romain came back from their first EVA they immediately left (without even taking their backpacks) to gather some geological samples for the SALM mission, following a request we received earlier in the morning to gather some soil samples at specific locations found yesterday by the parrot.

When they came back it was already 5pm but we had planned to visit another two sites to collect geological samples, placed at around 2 km and 4 km from the habitat, so I and Audrey departed on this mission hoping to be back before our infrared vision was needed.

After around 20 minutes we were already at the first location, a small cliff that was easy enough even with the spacesuit on so we decided to continue to the second location. Unfortunately, our navigation skills along with the Martian GPS equipment we have, led us on a big ATV marathon... but we did not quit. We really wanted those samples!

At around 6:00 pm we managed to arrive at the second site, and what a surprise... a big 15 meter cliff that we needed to climb to get the sample, and this is not so easy to do with our spacesuits on. We did not go all the way there to not get the sample so we climbed it, got the chance to see some Martian snow, and went back to the habitat with the Sun well under the horizon.

It was another perfect day on Mars.


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.76

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.54

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 11 Gallons


Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 11

· Trailer Potable Water Tank Level (inches from bottom): Full

· Hab Potable Water Tank Level (inches from bottom): 12

· Potable Water Meter Reading: 424386

Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: Green water

· Condition of plants in Tank 2: Green things

· Condition of plants in Tank 3: Green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little. As Joshua Nelson asked, I verified if the gray tank was still frozen. The grey tank is no longer frozen. So I used the pump to bring water in the straining tank. 2 leaks appeared on the pipe joints when I was pumping. 1 before the pump and 1 before the straining tank. I managed to fix them, screwing on the hose clamps.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:


Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car: just to supply in electricity the relay for the Rover during our first attempt to do mission SALM.


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: Yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status: Filled

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: Yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 70

· Thermostat setting downstairs (Farenheit): 60

· Actual temperature downstairs (Farenheit): 63


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:

As I said in my EVA report: we have a lot of problems with the radios and with the hands-free kits.


Science Report

Ayako Ono Reporting

1.     Habitability Experiment: This morning also, we started with habitability experiment. Ivo and Romain finished all their experiments, although some questionnaire and debriefing are needed.

2.     Ivo, Audrey, and Ayako worked on the EVA #10: The mission 2 on EVA #10 could not reach to the goal because of the technical issues on ATVs, although the first point was OK and safely came back.

3.     Ivo, Romain, and Kévin worked on the EVA #11: Mission 2 (testing the influence of stress related to communication on an EVA) was successfully done.

4.     Romain, and Kévin worked on the EVA #12: Collecting samples for the SALM mission. They successfully got samples from P2 P3.

5.     Ivo, and Audrey worked on the EVA #13: Collecting samples for ESTEC and SALM mission. They successfully got samples from P5 P6.

*Please see the details on their reports.

Thanks a lot for the suggestions to adjust colors for habitability experiments.


EVA 10 Report

Ayako Ono Reporting

EVA #10, 14:16-14:36

Crew Members: Ivo Ferreira, Audrey Lan San, Ayako Ono

Site Location: Waypoint 435, UTM 12S, 0519041 4251489; Waypoint 436, UTM 12S, 0519183 4251287

Transit Mode: ATVs
Objectives: Mission 2 (testing the influence of stress related to communication on an EVA).
Results: Had problems with ATVs several times. Managed 1 critical objective.

Lessons Learned: The person who can contact to mission control using GPS maybe better to save the time.


EVA 11 Report

Kévin Theatre Reporting

EVA #11, 15:00-15:40

Crew Members: Ivo Ferreira, Romain Benchenafi, Kévin Théatre

Site Location: foot: 0519041 4251489, 0519183 4251287, 0519477 4251280, 0519555 4251265

Transit Mode: ATVs

Objectives: Mission 2(testing the influence of stress related to communication on an EVA)

Results: Managed to comply with the 3 critical objectives.

Lessons Learned: Using bad radios is very penalizing in an EVA. Radios need to be change !


EVA 12 Report

Romain Benchenafi Reporting

EVA #12, 16:00 - 17:00

Crew members : THEATRE Kévin, BENCHENAFI Romain

Site location :

P2 : UTM 12S N38.40746 W110.79280

P3: UTM 12S N38.40737 W110.79261

Transit mode : Foot

Objective : Collect samples for ESTEC and SALM

Result : Samples retrieved

Lesson learned : It is very uncomfortable to collect samples with the EVA suits


EVA 13 Report

Audrey Lan San Reporting

EVA #13, 22th February 2011, 17:15 - 18:50

Crew members : Ivo Ferreira and Audrey Lan San.

Site location :

P5 : N 38.42638 W 110.78342

P6 : N 42.638 W 110.83420

Transit mode : ATV

Objective : Collecting samples for the SALM mission.

Result : We succeeded in reaching the locations of both samples we had to get.

Lesson learned : Stay on the routes as long as possible.


Chef's Report

Romain Benchenafi Reporting

1a) Today was  - a non-cooking day


1b) Was today a special day celebrated at the main meal with special food or activity? No.


2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.


Texas BBQ chicken with beans


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)


Nothing


4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.  None.


5)  Lessons learned  (if any)   No


6) Comments/questions for Kim and Jean, the food study investigators - N/A


Journalist Report

Romain Benchenafi Reporting

Today's afternoon was dedicated to EVAs. We are stillinvestigating for SALM, indeed we had been asked to collect samples from the sites P2 and P3, and we did so. It is a real pleasure to work in cooperation with this institute as it is the occasion for us to go further than our fields of study. Hope this will have a good conclusion!